The equation x^2 + x - 4 = (x + a) (x+ b) is true. what is ab
Answer by stanbon(75887)

The equation x^2 + x - 4 = (x + a) (x+ b) is true. what is ab
---
x^2+x-4 = x^2+(a+b)x+ab
----
Equations:
a+b = 1
ab = -4
--------------
Cheers,
Stan H.
==============